Use Laplace transforms to solve the following value problem

y''-y'=e^(t)cost y(0)=1 y'(0)=-1
Answer #1

Consider y-y = et cosct) Taking Laplace transform on both side, we we get 5 & {y? Lyn fet cas (t)} Bit s-1 To syl} = 52415Litres Ltd s3 – 452 +75-5 ? S(S-1) ((5 2 - 1/²+1) y(t) = et (sintt) - costly) - et + 7 This is required son of given I. v.P.